Find four consecutive even intergers whose sum is 340.
Answer by darq(90)

#1=x
#2=x+2
#3=x+4
#4=x+6

x+x+2+x+4+x+6=340
4x+12=340
4x=328
x=82

#1=82
#2=84
#3=86
#4=88